OpenStudy (anonymous):

How do i find the integral(or antiderivative) of (3-x)(x^2-6x)^5

myininaya (myininaya):

let u=the inside of nastier part.

myininaya (myininaya):

Hello?

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

u=x^2-6x du=2x-6=2(x-3)

myininaya (myininaya):

du=-2(3-x) dx -du/2=(3-x)dx

OpenStudy (anonymous):

What do I do with the 3-x though?

OpenStudy (anonymous):

replace (3-x)dx with -1/2 du

OpenStudy (anonymous):

Why do I replace it with -1/2 du?

OpenStudy (anonymous):

you want to convert the integral from integral(f(x)dx) to integral(f(u)du) if u=x^2-6x, then u'(x)=du/dx=2(x-3). thus du=dx u'(x)=2(x-3)dx thus -1/2*du=(3-x)dx the original integral is integrate( (3-x)(x^2-6x)^5 dx) =integrate( u^5 (3-x)dx) =integrate ( u^5*(-1/2) du) after integrating, convert back to x using the fact that u=x^2-6x

OpenStudy (anonymous):

OH! okay I understand now. Thank you!!